OpenStudy (anonymous):

Evaluate the function at the indicated value. Function: f(x)=x^3+7x Value: f(1/3)=

OpenStudy (pottersheep):

Just plug in 1/3 wherever you see x

OpenStudy (pottersheep):

f(1/3) = (1/3)^3 + 7(1/3) Can you solve that?

OpenStudy (pottersheep):

= (1/27) + (7/3)

OpenStudy (anonymous):

put x=1/3 in place of x

OpenStudy (pottersheep):

so I believe 64/27

OpenStudy (anonymous):

Thank you! I've been trying to figure out what I did wrong because webassign was being picky and wouldn't take the decimal form.

OpenStudy (pottersheep):

No problem :)

OpenStudy (igreen):

\(f(x) = \dfrac{1}{3}^3+7(\dfrac{1}{3})\) \(f(x) = \dfrac{1}{27}+\dfrac{7}{3}\) \(f(x) = 2\dfrac{10}{27}\)